# Confidence-weighted integration of human and machine judgments for superior decision-making

### v2

@@ Abstract
Large language models (LLMs) have emerged as powerful tools in various domains. Recent studies have shown that LLMs can surpass humans in certain tasks, such as predicting the outcomes of neuroscience studies. What role does this leave for humans in the overall decision process? One possibility is that humans, despite performing worse than LLMs, can still add value when teamed with them. A human and machine team can surpass each individual teammate when team members’ confidence is well-calibrated and team members diverge in which tasks they find difficult (i.e., calibration and diversity are needed). We simplified and extended a Bayesian approach to combining judgments using a logistic regression framework that integrates confidence-weighted judgments for any number of team members. Using this straightforward method, we demonstrated its effectiveness in both image classification and neuroscience forecasting tasks. Even when humans performed worse than machines, combining human judgments with one or more machines consistently improved overall team performance. Our hope is that this simple and effective strategy for integrating the judgments of humans and machines will lead to productive collaborations.
@@
Human-AI Collaboration, Large Language Models (LLMs), Neuroscience, Decision Making
@@ 1 Introduction
One potential solution is forming human-machine teams in which judgments from humans and machines are integrated [11, 12, 13]. It might be possible that humans can contribute to and make the overall team better even when their performance is worse on average than their machine teammates.
We will begin evaluating this possibility in an object recognition task where human and machine performance overlap according to experimental conditions [13]. Human-machine teaming combines the individual judgments of humans and machines. We will then evaluate a knowledge-intensive task in which large language models (LLMs) surpass humans in predicting the outcomes of neuroscience studies [14], posing a real challenge for effective team collaboration. Complementarity is realized when a team’s performance improves beyond that of either teammate alone [13, 12]. We investigate whether human-LLM teams outperform LLMs even when humans have inferior performance compared to LLMs. There are two key conditions for team complementarity to be fulfilled [14, 13, 15]. The first requirement is calibration of confidence. This implies that when humans and LLMs have a higher degree of confidence in their judgments, the accuracy of those judgments tends to be greater [14]. The second requirement is classification diversity among team members. Diversity holds when the errors in classification made by humans and LLMs are not the same [14].
Previous work [13] has explored the conditions for complementarity in the context of object recognition. Humans outperformed machines in the classification of natural images with low levels of noise, raising the question of whether a combined approach could achieve superhuman performance. They developed a Bayesian model that integrates the judgments of humans and machines. With this approach, human-machine complementarity was observed. However, the combination model is computationally expensive and challenging to extend to additional teammates. Ideally, a model that combines the judgments of humans and machines should be adaptable and scalable, easily interpretable, and allow for any number of teammates.
Here, we aim to offer this ideal solution to human-machine teaming while evaluating complementarity in an object recognition task, and a knowledge-intensive task that is not based on perceptual judgment. Critically, in both scenarios, humans were surpassed by machine systems. Foreshadowing our results, we find support for effective human-machine teaming through our novel and resource-efficient procedure. Our procedure comprises a logistic-regression-based strategy that provides confidence-weighted integration of teammates’ predictions for any number of team members. Our approach is particularly well-suited for combining human and machine judgments and assessing their contribution in predictive tasks.
@@ 2 Results
In this study, we explore whether humans can contribute to decisions when machine models, such as LLMs, are superior to them. We developed a logistic-regression-based method that integrates a weighted average of judgments from teammates, whether humans or machines. The proposed approach adheres to similar previously reported principles [13], offering a number of advantages: it is easy to use, flexible, and resource-efficient (details can be found in the Methods section). We evaluate our method in an object recognition task where effective human-machine collaboration has been demonstrated [13]. We then shift our focus to BrainBench [14] because LLMs significantly outperform human experts, posing a real challenge for team collaboration.
Figure 1: Performance of the confidence-weighted logistic combination model in a noisy object recognition task [13]. Accuracy results on high levels of image noise ($\Omega=125$) with the logistic combination model. Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model.
@@
We first assessed the performance of human-machine teams in the classification of noisy natural images [13]. The images were distorted by phase noise at each spatial frequency, where the phase noise was uniformly distributed in the interval $[-\Omega,\Omega]$ [16]. We considered two noise levels: images distorted by low ($\Omega=80$) and high ($\Omega=125$) noise. In the case of low noise, machines are surpassed by humans ($t(4)=-4.77$, $P<0.01$). The Bayesian combination model demonstrated human-machine complementarity (Figure S1). Despite its simpler setup, our confidence-weighted logistic combination model was able to provide team complementarity not only for human-machine teams but also for machine-machine teams (Figure S1). Furthermore, our approach compared favorably to the Bayesian combination model ($t_{Welch}(20)=2.91$, $P<0.01$).
@@
Figure 1 shows the performance of our confidence-weighted logistic combination model in the case of high noise, where most machines outperform humans. Of primary interest was whether teams including humans performed better than machine-only teams. We assessed machine-only teams comprising either one or two machines. Adding a human teammate to machine-only teams always improved the team’s performance ($t_{Welch}(22)=4.70$, $P<0.0001$).
@@
Figure 2: Assessing Humans and LLMs using BrainBench [14]. (a) The benchmark consisted of test cases constructed from the Journal of Neuroscience abstracts. Abstracts consist of background, methods, and results. The test-taker chose which of two versions of the abstract was the original version. The altered version maintained coherency while significantly altering the results. The 100 test cases considered here were constructed by GPT-4 with human oversight and quality control. (b) An example test case. Humans were instructed to select which version of the abstract was the original by clicking on either blue or green text to select that set of options. Test cases varied in the numbers of alternatives, but a single click will choose all options of the same color. After their choice, humans indicated their confidence. LLMs chose the version of the abstract that had the lower perplexity and their confidence was assessed by the absolute difference in perplexity of the two options.
@@
Figure 3: Conditions for effective collaboration between human experts and LLMs were satisfied. (a) When LLMs were confident in their BrainBench judgments, they were more likely to be correct. Confidence ratings were sorted into equal bins, and the mean accuracy for each bin was plotted. The positive slope of the black regression lines for Llama2 chat models (7B, 13B, and 70B) indicates well-calibrated confidence, meaning higher confidence correlated with higher accuracy. (b) LLMs surpassed human experts on BrainBench overall. Error bars represent standard error of the mean using a binomial model. (c) Item difficulty Spearman correlations among LLMs and human experts. For LLMs, difference in perplexity between incorrect and correct abstracts was used to determine the relative difficulty of test cases. Mean accuracy was used for human experts. LLMs align more with each other than humans, which implies human-machine teams will be diverse.
@@
Bayesian integration is marked by combining judgments based on confidence ratings from team members, whether human or machine. Our confidence model can use this information as well. However, in this dataset, is confidence weighting crucial or could improved team performance simply reflect an averaging or wisdom-of-the-crowd effect? This question has not been asked with this dataset, but is straightforward to evaluate using variations of our regression approach. We found that removing confidence in our approach did not negatively impact team performance in this noisy object recognition task (Figure S2). For human-machine teams, there is no difference between our standard model and a variant that does weight by confidence ($t_{Welch}(26)=0.29$, $P=0.39$). Surprisingly, for machine-machine teams, the no-confidence model outperformed the standard ($t_{Welch}(17)=2.58$, $P<0.01$). On the one hand, these results demonstrate the versatility of our approach in that different variations of our model can assess the basis for the success of human-machine teams. On the other hand, the fact that the object recognition dataset did not require confidence-weighted integration motivates considering another dataset for which we know that confidence of humans and machines are calibrated in that higher confidence is associated with higher accuracy. The next dataset considered provides a valuable test of our standard confidence-weighted model.
@@
Figure 4: Performance of all possible teams using the confidence-weighted logistic combination model. Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
@@
Calibration of confidence and classification diversity afford reliable human-LLM teaming
@@
We evaluated the conditions for effective collaboration (i.e., complementarity), namely calibration of confidence and classification diversity among team members. Both humans and LLMs were calibrated in that accuracy was positively correlated with confidence (Figure 3a). Diversity held in that LLMs and humans differed on which test items led to errors (Figure 3c). In terms of accuracy (Figure 3b), LLMs numerically surpassed humans ($t(2)=5.20$, $P<0.05$). Thus, we can consider whether humans can benefit teams consisting of machines that perform comparably or better.
@@
Similarly as in the image classification task, we investigated whether teams including humans performed better than LLM-only teams in the prediction of neuroscience results. All 15 possible team combinations, ranging from individual teammates to a 4-way human-LLM team, were considered (Figure 4). Adding a human teammate to LLM-only teams always improved the team’s performance ($t_{Welch}(8)=8.24$, $P<0.0001$).
@@
In the object recognition task, team performance is not affected by teammates’ confidence. Does the fluctuating confidence on a trial by trial basis matter for LLMs? Or could complementarity be achieved by just forming a weighted average of responses as in the object recognition task? We evaluated the impact of confidence by setting the magnitude of the confidence scores to 1. This setting mimics the concept of “the wisdom of the crowds,” where responses are considered without factoring in confidence levels. We find that confidence is particularly important in this knowledge-intensive task, as the overall performance diminishes when confidence scores are neglected (Figure 5). With this variant, human-LLM teams do not always surpass LLM-only teams ($t_{Welch}(7)=0.38$, $P=0.36$). The model including confidence (Figure 4) outperforms the variant without confidence (Figure 5) by a significant margin ($t_{Welch}(17)=3.16$, $P<0.01$).
@@
We then assessed whether our regression approach would benefit from a more elaborated formulation. Eq. (3) presents a function that modulates the magnitude of a team member’s confidence rating to adjust their calibration. In an optimal setting (Figure S3), the resulting model was indistinguishable from the standard regression model ($t_{Welch}(19)=0.05$, $P=0.48$). Adding interaction terms to the confidence-weighted features does not improve the performance of either human-LLM or LLM-only teams (Figure S4). More complex model variants may perform better in other tasks, especially when there is more and less noisy data. Simpler model variants seem to be more robust to noise in our cross-validated experiments. Thus, the simplicity of our approach provides effective team collaboration.
@@
Figure 5: Removing confidence from the logistic combination model diminishes team performance. Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where confidence scores were set to 1, i.e., $f(x)=1$ in Eq. (3). Adding a human to a team with one or more machines (blue points) does not necessarily improve performance (green points). Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
@@
Our confidence-weighted logistic regression approach follows from the principles of a Bayesian combination model that fosters human-machine complementarity [13]. One questions is how well our logistic regression approach compares to the Bayesian approach beyond image classification. Our confidence-weighted regression model compared favorably (Figure 6) to the Bayesian model when evaluated on the three human-LLM teams for which the Bayesian model is intended to apply ($t_{Welch}(2)=4.08$, $P<0.05$). This success is impressive given that the regression approach takes seconds to compute on a current desktop whereas the Bayesian approach would take 12-13 days.
@@ 3 Discussion
Can humans team effectively with machines such as LLMs when the humans perform worse? We developed a confidence-weighted regression approach that can integrate judgments from any number of teammates. Using this method and testing on two forecasting benchmarks [13, 14], we found that human-machine teams achieve complementarity, that is their combined performance bests that of either teammate alone (Figures 1 and 4). Complementarity was achieved because two critical conditions were satisfied, namely confidence was well-calibrated and classification diversity held among teammates (Figure 3). Strikingly, every combination of machines benefited from adding a human to the team (Figures 1 and 4).
Figure 6: Comparison between Bayesian and confidence-weighted logistic combination models for human-LLM teams. Each data point corresponds to the average across 503 test case evaluations. Llama2 chat 7B, 13B, and 70B models were considered. Error bars represent standard error of the mean using a binomial model. The confidence-weighted logistic combination model more effectively integrated human and machine judgments.
Confidence in individual responses significantly impacted team performance in the knowledge-intensive task. The overall performance diminished when confidence scores were neglected (Figure 5). However, in the noisy object recognition task, this effect was not present (Figure S2). In both scenarios, the performance of the confidence-weighted integration of judgments from teammates was equal to or better than the weighted average of responses, i.e., “the wisdom of the crowds.” For the confidence-weighted model variants to thrive, confidence judgments need to be calibrated such that higher confidence is assocaited with higher accuracy.
@@
While we selected three LLMs with superhuman performance on BrainBench, these LLMs are not the highest performing models on this benchmark [14]. Our choice was deliberate because a vastly superior teammate may hinder complementarity. In the limit, a teammate who is never wrong does not need to be part of a team. This limiting condition may become more prevalent should LLMs continue to improve and, therefore, diminish the benefits of human-LLM teaming. For the foreseeable future, we suspect there will be tasks for which humans and LLMs can effectively team. Moreover, our method for integrating the judgments of teammates is not limited to human-LLM teams. Instead, the method is general and applies to any set of agents (natural or artificial) that can report how confident they are in their decisions.

### v3

@@ Abstract
Large language models (LLMs) can surpass humans in certain forecasting tasks. What role does this leave for humans in the overall decision process? One possibility is that humans, despite performing worse than LLMs, can still add value when teamed with them. A human and machine team can surpass each individual teammate when team members’ confidence is well-calibrated and team members diverge in which tasks they find difficult (i.e., calibration and diversity are needed). We simplified and extended a Bayesian approach to combining judgments using a logistic regression framework that integrates confidence-weighted judgments for any number of team members. Using this straightforward method, we demonstrated its effectiveness in both image classification and neuroscience forecasting tasks. Combining human judgments with one or more machines consistently improved overall team performance. Our hope is that this simple and effective strategy for integrating the judgments of humans and machines will lead to productive collaborations.
@@
Human-AI Collaboration, Teaming, Complementarity, Large Language Models (LLMs), Object Recognition, Neuroscience, Decision Making
@@ 1 Introduction
One potential solution is forming human–machine teams in which judgments from humans and machines are integrated [11, 12, 13]. It might be possible that humans can contribute to and make the overall team better even when their performance is worse on average than their machine teammates.
We will begin evaluating this possibility in an object recognition task where human and machine performance overlap according to experimental conditions [13]. Human–machine teaming combines the individual judgments of humans and machines. We will then evaluate a knowledge-intensive task in which large language models (LLMs) surpass humans in predicting the outcomes of neuroscience studies [14], posing a real challenge for effective team collaboration. Complementarity is realized when a team’s performance improves beyond that of either teammate alone [12, 13]. We investigate whether human–LLM teams outperform LLMs even when humans have inferior performance compared to LLMs. There are two key conditions for team complementarity to be fulfilled [13, 14, 15]. The first requirement is calibration of confidence. This implies that when humans and LLMs have a higher degree of confidence in their judgments, the accuracy of those judgments tends to be greater [14]. The second requirement is classification diversity among team members. Diversity holds when the errors in classification made by humans and LLMs are not the same [14].
Previous work [13] has explored the conditions for complementarity in the context of object recognition. Humans outperformed machines in the classification of natural images with low levels of noise, raising the question of whether a combined approach could achieve superhuman performance. They developed a Bayesian model that integrates the judgments of humans and machines. With this approach, human–machine complementarity was observed. However, the combination model is computationally expensive and challenging to extend to additional teammates. Ideally, a model that combines the judgments of humans and machines should be adaptable and scalable, easily interpretable, and allow for any number of teammates.
Here, we aim to offer this ideal solution to human–machine teaming while evaluating complementarity in an object recognition task, and a knowledge-intensive task that is not based on perceptual judgment. Critically, in both scenarios, humans were surpassed by machine systems. Foreshadowing our results, we find support for effective human–machine teaming through our resource-efficient procedure. Our procedure comprises a logistic-regression-based strategy that provides confidence-weighted integration of teammates’ predictions for any number of team members. Our approach is particularly well-suited for combining human and machine judgments and assessing their contribution in predictive tasks.
@@ 2 Results
In this study, we explore whether humans can contribute to decisions when machine models, such as LLMs, are superior to them. We developed a logistic-regression-based method that integrates a weighted average of judgments from teammates, whether humans or machines. The proposed approach adheres to similar previously reported principles [13], offering a number of advantages: it is easy to use, flexible, and resource-efficient (details can be found in the Methods section). We evaluate our method in an object recognition task where effective human–machine collaboration has been demonstrated [13]. We then shift our focus to BrainBench [14] because LLMs significantly outperform human experts, posing a real challenge for team collaboration.
Figure 1: Performance of the confidence-weighted logistic combination model in the noisy object recognition task [13].
Accuracy results on high levels of image noise ($\Omega=125$) with the logistic combination model. Human–machine teams (green points) consistently outperform teams without humans (blue points). Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model.
@@
We first assessed the performance of human–machine teams in the classification of noisy natural images [13]. The images were distorted by phase noise at each spatial frequency, where the phase noise was uniformly distributed in the interval $\Omega,\Omega] [$16]. We considered two noise levels: images distorted by low ($\Omega=80$) and high ($\Omega=125$) noise. In the case of low noise, machines are surpassed by humans (t(4)=4.77t(4)=-4.77, $P<0.01$). The Bayesian combination model [13] demonstrated human–machine complementarity (Figure S1). Despite its simpler setup, our confidence-weighted logistic combination model was able to provide team complementarity not only for human–machine teams but also for machine–machine teams (Figure S2). Furthermore, our approach outperformed the Bayesian combination model (Welch’s $t(20.87)=2.91$, $P<0.01$).
@@
Figure 1 shows the performance of our confidence-weighted logistic combination model in the case of high noise, where most machines outperform humans (Figure S3). Of primary interest was whether teams including humans performed better than machine-only teams. We assessed machine-only teams comprising either one or two machines. Adding a human teammate to machine-only teams always improved the team’s performance (Welch’s $t(22.78)=4.70$, $P<0.0001$).
Figure 2: Assessing Humans and LLMs using BrainBench [14].
@@
(A) The benchmark comprises test cases constructed from the Journal of Neuroscience abstracts. Abstracts consist of background, methods, and results. The test-taker chose which of two versions of the abstract was the original version. The altered version maintained coherency while significantly altering the results. The 100 test cases considered here were constructed by GPT-4 with human oversight and quality control.
(B) An example test case. Humans were instructed to select which version of the abstract was the original by clicking on either blue or green text to select that set of options. Test cases varied in the numbers of alternatives, but a single click will choose all options of the same color. After their choice, humans indicated their confidence. LLMs chose the version of the abstract that had the lower perplexity score and their confidence was assessed by the absolute difference in perplexity of the two options.
Figure 3: Conditions for effective collaboration between human experts and LLMs were satisfied.
@@
(A) When human experts and LLMs were confident in their BrainBench judgments, they were more likely to be correct. Confidence ratings were sorted into equal bins, and the mean accuracy for each bin was plotted. The positive slope of the black regression lines for humans and Llama2 chat models (7B, 13B, and 70B) indicates well-calibrated confidence [14, 17, 18, 19], meaning higher confidence correlates with higher accuracy.
(B) Item difficulty Spearman correlations among LLMs and human experts. For LLMs, difference in perplexity between incorrect and correct abstracts was used to determine the relative difficulty of test cases. Mean accuracy was used for human experts. LLMs align more with each other than humans, which implies human–machine teams will be diverse. Heatmap color scale ranges from 0.1 to 0.9.
(C) LLMs surpass human experts on BrainBench overall. Error bars represent standard error of the mean using a binomial model.
@@
Bayesian integration is marked by combining judgments based on confidence ratings from team members, whether human or machine. Our confidence model can use this information as well. However, it remains unclear whether confidence weighting is crucial, or if improved team performance simply reflects an averaging or wisdom-of-the-crowd effect. This question has not been asked with this dataset, but is straightforward to evaluate using variations of our regression approach. We found that removing confidence in our approach did not negatively impact team performance in this noisy object recognition task (Figure S4). For human–machine teams, there is no difference between our standard model and the variant that does not weight by confidence (Welch’s $t(26.71)=0.29$, $P=0.77$). Surprisingly, for machine–machine teams, the no-confidence model outperformed the standard (Welch’s $t(17.47)=2.58$, $P<0.01$). Similar performance between signed confidence and sign-only features indicates that confidence does not contribute to improved predictions. This is a consequence of machine classifiers being skewed toward high probability scores (Figure S5). On the one hand, these results demonstrate the versatility of our approach in that different variations of our model can assess the basis for the success of human–machine teams. On the other hand, the fact that the object recognition dataset did not require confidence-weighted integration motivates considering another dataset for which we know that confidence of humans and machines are calibrated in that higher confidence is associated with higher accuracy. The next dataset considered provides a valuable test of our standard confidence-weighted model.
Figure 4: Performance of all possible teams using the confidence-weighted logistic combination model.
@@
Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Llama2 chat 7B, 13B, and 70B models are considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
@@
Calibration of confidence and classification diversity afford reliable human–LLM teaming
@@
We evaluated the conditions for effective collaboration (i.e., complementarity), namely calibration of confidence and classification diversity among team members. Both humans and LLMs were calibrated in that accuracy was positively correlated with confidence (Figure 3A). Diversity held in that LLMs and humans differed on which test items led to errors (Figure 3B). In terms of accuracy (Figure 3C), LLMs numerically surpassed humans by a small margin ($t(2)=5.20$, $P<0.05$). Thus, we can consider whether humans can benefit teams consisting of machines that perform comparably or better.
@@
Similarly as in the image classification task, we investigated whether teams including humans performed better than LLM-only teams in the prediction of neuroscience results. All 15 possible team combinations, ranging from individual teammates to a 4-way human–LLM team, were considered (Figure 4). Adding a human teammate to LLM-only teams always improved the team’s performance (Welch’s $t(8.29)=8.24$, $P<0.0001$). Pairing a human with an LLM led to a more effective (i.e., accurate) team than pairing the LLM with a different LLM ($t(2)=10.39$, $P<0.01$).
@@
In the object recognition task, team performance is not affected by teammates’ confidence. Does the fluctuating confidence on a trial by trial basis matter for LLMs? Or could complementarity be achieved by just forming a weighted average of responses as in the object recognition task? We evaluated the impact of confidence by setting the magnitude of the confidence scores to 1. This setting mimics the concept of “the wisdom of the crowds,” where responses are considered without factoring in confidence levels. We find that confidence is particularly important in this knowledge-intensive task, as the overall performance diminishes when confidence scores are neglected (Figure 5). With this variant, human–LLM teams do not always surpass LLM-only teams (Welch’s $t(7.36)=0.38$, $P=0.36$). The model including confidence (Figure 4) outperforms the variant without confidence (Figure 5) by a significant margin (Welch’s $t(17.08)=3.16$, $P<0.01$).
@@
We then assessed whether our regression approach would benefit from a more elaborated formulation. Equation (4) presents a function that modulates the magnitude of a team member’s confidence rating to adjust their calibration. In an optimal setting (Figure S6), the resulting model was indistinguishable from the standard regression model (Welch’s $t(19.32)=0.05$, $P=0.96$). Adding interaction terms to the confidence-weighted features does not improve the performance of either human–LLM or LLM-only teams (Figure S7). More complex model variants may perform better in other tasks, especially when there is more and less noisy data. Simpler model variants seem to be more robust to noise in our cross-validated experiments. Thus, the simplicity of our approach provides effective team collaboration.
Figure 5: Removing confidence from the logistic combination model diminishes team performance.
@@
Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where the magnitude of the confidence scores was set to 1, i.e., $f(x)=1$ in Equation (4). Adding a human to a team with one or more machines (blue points) does not necessarily improve performance (green points). Llama2 chat 7B, 13B, and 70B models are considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
@@
Our confidence-weighted logistic regression approach follows from the principles of a Bayesian combination model that fosters human–machine complementarity [13] (Figure S1). One questions is how well our logistic regression approach compares to the Bayesian approach beyond image classification. Our confidence-weighted regression model outperformed (Figure 6) the Bayesian model when evaluated on the three human–LLM and three LLM–LLM teams for which the Bayesian model is intended to apply (Welch’s $t(8.75)=2.91$, $P<0.01$). This success is impressive given that the regression approach takes seconds to compute on a current desktop, whereas the Bayesian approach is orders of magnitude slower.
@@ 3 Discussion
Can humans team effectively with machines such as LLMs when the humans perform worse? We developed a confidence-weighted regression approach that can integrate judgments from any number of teammates. Using this method and testing on two forecasting benchmarks [13, 14], we found that human–machine teams achieve complementarity, that is their combined performance bests that of either teammate alone (Figures 1 and 4). Complementarity was achieved because two critical conditions were satisfied, namely confidence was well-calibrated and classification diversity held among teammates (Figure 3). Strikingly, every combination of machines benefited from adding a human to the team (Figures 1 and 4).
Figure 6: Comparison between Bayesian and confidence-weighted logistic combination models for human–LLM and LLM–LLM teams.
The confidence-weighted logistic combination model more effectively integrates human and machine judgments. Llama2 chat 7B, 13B, and 70B models are considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
Confidence in individual responses significantly impacted team performance in the knowledge-intensive task. The overall performance diminished when confidence scores were neglected (Figure 5). However, in the noisy object recognition task, this effect was not present (Figure S4). In both scenarios, the performance of the confidence-weighted integration of judgments from teammates was equal to or better than the weighted average of responses, i.e., “the wisdom of the crowds.” For the confidence-weighted model variants to thrive, confidence judgments need to be calibrated such that higher confidence is associated with higher accuracy.
@@
While we selected three LLMs with superhuman performance on BrainBench, these LLMs are not the highest performing models on this benchmark [14]. Our choice was deliberate because a vastly superior teammate may hinder complementarity. In the limit, a teammate who is never wrong does not need to be part of a team. This limiting condition may become more prevalent should LLMs continue to improve and, therefore, diminish the benefits of human–LLM teaming. For the foreseeable future, we suspect there will be tasks for which humans and LLMs can effectively team. Moreover, our method for integrating the judgments of teammates is not limited to human–LLM teams. Instead, the method is general and applies to any set of agents (natural or artificial) that can report how confident they are in their decisions.